# ID-centric Pre-training for Recommendation

### v1

@@ 3. Experiments
In this section, we aim to answer the following questions: (1) RQ1:How does IDP perform compared with other state-of-the-art (SOTA) baselines? (2) RQ2: Can our proposed framework be generalized to other base models? Furthermore, can our IDP be generalized to other downstream sequential models? (3) RQ3: How does IDP perform in the zero-shot setting? (4) RQ4: What are the effects of different modules in IDP? (5) RQ5: Can our IDP be utilized for helping inner-domain recommendation? (6) RQ6: Can we derive some interesting insights by analyzing IDPs?
@@ 3.1. Dataset
Following (He et al., 2017), we transformed all rating datasets into the implicit dataset, in which each instance are labeled with 0 or 1 to indicate whether a user rated an item. whether the user has rated the item For the pre-training datasets, following previous works (Kang and McAuley, 2018; Zhou et al., 2020), we discard the users and items with less than five interactions, as the cold start problem of the pre-training domain is not our concern. We first conduct pre-training on all five pre-trained models. Subsequently, we fine-tune the pre-trained model separately on five downstream datasets to validate the proposed method. For all datasets, We adopt the title of products as the description text of items.
@@
Dataset Metric GRU4Rec SASRec BERT4Rec DACDR PterRec FDSA Recformer UniSRecID+t IDP IDPID+t Impro.
Arts NDCG@1 0.2343 0.3228 0.2446 0.2666 0.2713 0.2996 0.3281 0.3841 0.3404 0.4176 8.02%
NDCG@3 0.3346 0.4027 0.3485 0.3601 0.3613 0.4005 0.4137 0.4926 0.4249 0.5153 4.41%
NDCG@5 0.3698 0.4298 0.3837 0.3923 0.3905 0.4341 0.4438 0.5261 0.4528 0.5454 3.54%
HR@3 0.4071 0.4602 0.4234 0.4275 0.4259 0.4731 0.4754 0.5699 0.4855 0.5849 2.56%
HR@5 0.4928 0.5262 0.5091 0.5059 0.4968 0.5546 0.5486 0.6514 0.5532 0.6581 1.02%
MRR 0.3569 0.4211 0.3691 0.3805 0.3799 0.4219 0.4366 0.5081 0.4423 0.5302 4.17%
Instru- ments NDCG@1 0.2382 0.3121 0.2623 0.2774 0.2485 0.3042 0.2812 0.3404 0.3286 0.3977 14.41%
NDCG@3 0.3361 0.3959 0.3648 0.3858 0.3409 0.4108 0.3382 0.4492 0.4207 0.4990 9.98%
NDCG@5 0.3728 0.4243 0.4017 0.4219 0.3747 0.4463 0.3719 0.4846 0.4504 0.5305 8.65%
HR@3 0.4073 0.4562 0.4390 0.4640 0.4077 0.4873 0.4015 0.5274 0.4865 0.5715 7.72%
HR@5 0.4964 0.5252 0.5286 0.5517 0.4901 0.5737 0.4835 0.6135 0.5588 0.6481 5.34%
MRR 0.3618 0.4149 0.3877 0.4047 0.3655 0.4318 0.3666 0.4686 0.4388 0.5149 8.99%
Office NDCG@1 0.2884 0.3408 0.3101 0.2879 0.3082 0.3317 0.2521 0.3899 0.3458 0.4096 4.81%
NDCG@3 0.4052 0.4429 0.4296 0.4054 0.4155 0.4458 0.3142 0.5055 0.4487 0.5186 2.53%
NDCG@5 0.4460 0.4790 0.4681 0.4462 0.4501 0.4832 0.3382 0.5418 0.4831 0.5528 1.99%
HR@3 0.4896 0.5169 0.5155 0.4907 0.4927 0.5278 0.3590 0.5882 0.5227 0.5965 1.39%
HR@5 0.5887 0.6047 0.6089 0.5896 0.5768 0.6185 0.4176 0.6763 0.6088 0.6798 0.51%
MRR 0.4267 0.4619 0.4467 0.4245 0.4329 0.4645 0.3429 0.5205 0.4668 0.5331 2.36%
Prime NDCG@1 0.1094 0.1520 0.0950 0.1192 0.1119 0.1209 0.1519 0.1280 0.1526 0.1842 21.18%
NDCG@3 0.1822 0.2224 0.1561 0.1896 0.1804 0.1877 0.2069 0.1991 0.2268 0.2630 16.99%
NDCG@5 0.2163 0.2536 0.1861 0.2214 0.2136 0.2188 0.2298 0.2313 0.2595 0.2953 16.44%
HR@3 0.2361 0.2743 0.2013 0.2415 0.2314 0.2370 0.2450 0.2518 0.2804 0.3209 16.99%
HR@5 0.3193 0.3503 0.2747 0.3191 0.3123 0.3127 0.3010 0.3302 0.3598 0.3996 14.07%
MRR 0.2219 0.2584 0.1977 0.2268 0.2209 0.2261 0.2395 0.2372 0.2629 0.2965 16.37%
@@ 3.2. Experimental Setting
Evaluation Protocols. We adopt the leave-one-out strategy to evaluate the models’ performance. Following (Kang and McAuley, 2018; Zhou et al., 2020), for each ground truth, we randomly sample $99$ items that the user did not interact with under the target behavior as negative samples. Besides, for a more accurate comparison, we also conducted experiments on the setting using full candidates. As this setting is too time-consuming. we compare the best baseline with our IDP. The results show IDP still demonstrated significant improvements across all metrics. We employ top-K hit rate (HR@k), top-K (NDCG@k), and MRR, with $k=\{1,3,5\}$.
@@ 3.4. Overall Performance Comparison (RQ1)
The main results of our IDP are shown in Table 2. We have:
@@ 3.5.3. The universality of IDP on the cross-platform scenarios
Table 5. Results on the cross-platform scenario. All improvements are significant (t-test with p <{<}< 0.05).
Dataset Metric GRU4Rec SASRec BERT4Rec DACDR PterRec FDSA Recformer UniSRecID+t IDP IDPID+t Impro.
ML-1M NDCG@1 0.3374 0.3566 0.3258 0.1444 0.2187 0.3313 0.1332 0.2902 0.3606 0.3566 1.10%
NDCG@3 0.4794 0.4948 0.4668 0.2278 0.3475 0.4667 0.2256 0.4328 0.4966 0.4904 0.36%
NDCG@5 0.5253 0.5362 0.5117 0.2649 0.3961 0.5107 0.2685 0.4810 0.5381 0.5363 0.35%
HR@3 0.5808 0.5927 0.5682 0.2891 0.4417 0.5644 0.2942 0.5361 0.5935 0.5874 0.10%
HR@5 0.6919 0.6930 0.6765 0.3791 0.5599 0.6710 0.3987 0.6531 0.6942 0.6987 0.80%
MRR 0.4945 0.5082 0.4821 0.2675 0.3763 0.4840 0.2629 0.4513 0.5105 0.5062 0.45%
@@ 3.8. Representing Intra-domain Cold-start Items via Pre-trained Item Embeddings (RQ5)
In this section, we aim to investigate the effects of generating IDs within the inner domain. This experimental result has the ability to demonstrate whether our IDP is capable of addressing inner-domain cold start issues. Instead of generating ID embeddings from the upstream domain to the downstream domain, we focus on generating ID embeddings within the inner domain. Specifically, for a domain $A$, we first train a sequential model. Then we use our IDP to generalize ID embedding for cold items (without any interaction). Last we directly evaluate the model’s performance on cold items without any tuning operation. Here, we randomly select 10% items as cold items and remove them from the training dataset. The results are shown in Table LABEL:tab:inner_domain, where IDPInner is the model that re-generates ID embedding by our CDIM in the Inner domain. From this table, we can observe that our IDP can still provide reasonably accurate recommendations. This indicates that our IDP can be flexibly applied to inner-domain item cold start scenarios.
@@
Dataset Metric HR@5 NDCG@5 HR@10 NDCG@10
Instruments IDPInner 0.2352 0.1819 0.3110 0.2063
Arts IDPInner 0.2705 0.1874 0.4203 0.2355
@@ 3.9. Analysis of Sequential Encoder (RQ6)
Dataset Metric HR@5 NDCG@5 HR@10 NDCG@10
Instruments IDP 0.5588 0.4504 0.6532 0.4810
IDP w/o PS 0.5522 0.4500 0.6379 0.4778
Prime IDP 0.3598 0.2595 0.4865 0.3002
IDP w/o PS 0.3715 0.2693 0.4996 0.3106
@@
Dataset Metric HR@5 NDCG@5 HR@10 NDCG@10
Instruments SASRec 0.5252 0.4243 0.6162 0.4538
SASRecEX 0.4903 0.3846 0.5842 0.4151
Prime SASRec 0.3503 0.2536 0.4771 0.2944
SASRecEX 0.2946 0.2092 0.4103 0.2464
4. Related Works
Sequential Recommendation. Sequential recommendation models mainly leverage users’ chronological behavior sequences to learn users’ preferences. Recently, with the great progress of deep learning, ID-based sequential models have achieved great success, which project item ID to unique dense embedding. GRU4Rec (Hidasi et al., 2016) proposes to use Gated Recurrent Units in the session-based recommendation. Inspired by the success of Transformer and BERT (Vaswani et al., 2017; Devlin et al., 2019), SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) adopt self-attention mechanisms to model user behavior sequences.
Pre-training in Recommendation. Recently, pre-training models have achieved great success in NLP (Devlin et al., 2019; Liu et al., 2019) and CV (He et al., 2021; Radford et al., 2021). They aim to transfer prior knowledge from general large-scale datasets to downstream domains. While in recommendation, since IDs are unaligned, the well-learned unique ID-embeddings are challenging to transfer directly to downstream new domains. To utilize the power of pre-trained model, some single-domain pre-trained models (Yao et al., 2021; Wu et al., 2021) have been proposed to improve recommendation accuracy of the inner domain. However, these models lack transferability and are unable to apply to cross-domain recommendation. Some works transfer well-learned knowledge by overlapped users/items (Hu et al., 2018; Tsai, 2018; Zhu et al., [n. d.], 2022; Zang et al., 2022; Yuan et al., 2020b). While overlapped users/items generally are limited in the real world, thus they are lack of universality. To build a universal model, some works attempt to design a text-based/modality-based pre-trained model (Hou et al., 2022; Ding et al., 2021; Cui et al., 2022). Those works introduce pre-trained language models (PLM) into recommendations and attempt to replace ID-embedding with modality information. Specifically, UniSRec (Hou et al., 2022) utilizes PLM to learn universal item representations across different domains via items’ textual information. However, recent studies show that ID-embedding is still irreplaceable (Gao et al., 2023; Liu et al., 2023; Hou et al., 2023). Different from previous works, we build an ID-centric pre-trained model, which directly transfers pre-trained ID-embedding to downstream domains.

### v2

@@ 4. Experiments
In this section, we aim to answer the following questions:
  • RQ1:How does IDP perform compared with other state-of-the-art (SOTA) baselines?
  • RQ2: Can our proposed framework be generalized to other base models? Furthermore, can our IDP be generalized to other downstream sequential models?
  • RQ3: How does IDP perform in the zero-shot setting?
  • RQ4: What are the effects of different modules in IDP?
  • RQ5: Can our IDP be utilized for helping inner-domain recommendation?
  • RQ6: Can our IDP be extended to multi-modal information?
  • RQ7: Can we derive some interesting insights by analyzing IDPs?
@@ 4.1. Dataset
Following (He et al., 2017b), we transformed all rating datasets into the implicit dataset, in which each instance are labeled with 0 or 1 to indicate whether a user rated an item. whether the user has rated the item For the pre-training datasets, following previous works (Kang and McAuley, 2018; Zhou et al., 2020), we discard the users and items with less than five interactions, as the cold start problem of the pre-training domain is not our concern. We first conduct pre-training on all five pre-trained models. Subsequently, we fine-tune the pre-trained model separately on five downstream datasets to validate the proposed method. For all datasets, We adopt the title of products as the description text of items.
@@
Dataset Model NDCG@1 NDCG@3 NDCG@5 HR@3 HR@5 MRR
Arts GRU4Rec 0.2343 0.3346 0.3698 0.4071 0.4928 0.3569
SASRec 0.3228 0.4027 0.4298 0.4602 0.5262 0.4211
Bert4Rec 0.2446 0.3485 0.3837 0.4234 0.5091 0.3691
DACDR 0.2666 0.3601 0.3923 0.4275 0.5059 0.3805
PterRec 0.2713 0.3613 0.3905 0.4259 0.4968 0.3799
FDSA 0.2996 0.4005 0.4341 0.4731 0.5546 0.4219
Recformer 0.3281 0.4137 0.4438 0.4754 0.5486 0.4366
UniSrecID+t 0.3841 0.4926 0.5261 0.5699 0.6514 0.5081
IDP 0.3404 0.4249 0.4528 0.4855 0.5532 0.4423
IDPID+t 0.4176 0.5153 0.5454 0.5849 0.6581 0.5302
Impro 8.02% 4.41% 3.54% 2.56% 1.02% 4.17%
Instruments GRU4Rec 0.2382 0.3361 0.3728 0.4073 0.4964 0.3618
SASRec 0.3121 0.3959 0.4243 0.4562 0.5252 0.4149
Bert4Rec 0.2623 0.3648 0.4017 0.4390 0.5286 0.3877
DACDR 0.2774 0.3858 0.4219 0.4640 0.5517 0.4047
PterRec 0.2485 0.3409 0.3747 0.4077 0.4901 0.3655
FDSA 0.3042 0.4108 0.4463 0.4873 0.5737 0.4318
Recformer 0.2812 0.3382 0.3719 0.4015 0.4835 0.3666
UniSRecID+t 0.3404 0.4492 0.4846 0.5274 0.6135 0.4686
IDP 0.3286 0.4207 0.4504 0.4865 0.5588 0.4388
IDPID+t 0.3977 0.4990 0.5305 0.5715 0.6481 0.5149
Impro. 14.41% 9.98% 8.65% 7.72% 5.34% 8.99%
Office GRU4Rec 0.2884 0.4052 0.4460 0.4896 0.5887 0.4267
SASRec 0.3408 0.4429 0.4790 0.5169 0.6047 0.4619
Bert4Rec 0.3101 0.4296 0.4681 0.5155 0.6089 0.4467
DACDR 0.2879 0.4054 0.4462 0.4907 0.5896 0.4245
PterRec 0.3082 0.4155 0.4501 0.4927 0.5768 0.4329
FDSA 0.3317 0.4458 0.4832 0.5278 0.6185 0.4645
Recformer 0.2521 0.3142 0.3382 0.3590 0.4176 0.3429
UniSRecID+t 0.3899 0.5055 0.5418 0.5882 0.6763 0.5205
IDP 0.3458 0.4487 0.4831 0.5227 0.6088 0.4668
IDPID+t 0.4096 0.5186 0.5528 0.5965 0.6798 0.5331
Impro 4.81% 2.53% 1.99% 1.39% 0.51% 2.36%
Prime GRU4Rec 0.1094 0.1822 0.2163 0.2361 0.3193 0.2219
SASRec 0.1520 0.2224 0.2536 0.2743 0.3503 0.2584
Bert4Rec 0.0950 0.1561 0.1861 0.2013 0.2747 0.1977
DACDR 0.1192 0.1896 0.2214 0.2415 0.3191 0.2268
PterRec 0.1119 0.1804 0.2136 0.2314 0.3123 0.2209
FDSA 0.1209 0.1877 0.2188 0.2370 0.3127 0.2261
Recformer 0.1519 0.2069 0.2298 0.2450 0.3010 0.2395
UniSRecID+t 0.1280 0.1991 0.2313 0.2518 0.3302 0.2372
IDP 0.1526 0.2268 0.2595 0.2804 0.3598 0.2629
IDPID+t 0.1842 0.2630 0.2953 0.3209 0.3996 0.2965
Impro. 21.18% 16.99% 16.44% 16.99% 14.07% 16.37%
@@ 4.2. Experimental Setting
Evaluation Protocols. We adopt the leave-one-out strategy to evaluate the models’ performance. Following (Kang and McAuley, 2018; Zhou et al., 2020), for each ground truth, we randomly sample $99$ items that the user did not interact with under the target behavior as negative samples. We employ top-K hit rate (HR@K), top-K normalized discounted cumulative gain (NDCG@K), and mean Reciprocal Rank ( MRR), with $k=\{1,3,5\}$.
Specifically, the HR@k is calculated by the following:
(20) $\begin{split}HR@k=\frac{1}{|U|}\sum_{u\in U}\delta(\hat{R}(u,k)\cap R(u)\neq\emptyset),\end{split}$
where $\hat{R}(u,k)$ is the top-k predicted item list of user $u$ that sorting by predicted score. $R(u)$ is a ground-truth set of items of user $u$. $\delta(\cdot)$ is the indicator function, where the $\delta(b)=1$ if $b$ is true and 0 if false.
The $NDCG@k$ calculated by the following:
(21) $\qquad\operatorname{rel}(u,i)=\delta({\hat{R}(u,k)}^{i}\in R(u)),$
$\operatorname{NDCG}(N)$ $=\frac{1}{Y}\sum_{i=1}^{N}\frac{2^{\operatorname{rel}(u,i)}-1}{\log(i+1)}\text{ where }Y=\sum_{i=1}^{N}\frac{1}{\log(i+1)},$
where ${\hat{R}(u,k)}^{i}$ is the $i$-th item of ranked item list $\hat{R}(u,k)$. $\delta(\cdot)$ is the indicator function.
The $MRR$ is calculated by the following:
(22) $\begin{split}MRR=\frac{1}{|U|}\sum_{u\in U}rank_{u}\end{split}$
$rank_{u}$ is the rank position of the first relevant item in $\hat{R(u)}$.
@@ 4.4. Overall Performance Comparison (RQ1)
The main results of our IDP are shown in Table LABEL:tab:main_table. We have:
@@ 4.5.3. The universality of IDP on the cross-platform scenarios
Table 6. Results on the cross-platform scenario. The best and the second-best performances are bold and underlined. ”Impro.” denotes the relative improvement over the best baselines. All improvements are significant (t-test with p <{<}< 0.05).
Metric NDCG@1 NDCG@3 NDCG@5 HR@3 HR@5 MRR
GRU4Rec 0.3374 0.4794 0.5253 0.5808 0.6919 0.4945
SASRec 0.3566 0.4948 0.5362 0.5927 0.6930 0.5082
BERT4Rec 0.3258 0.4668 0.5117 0.5682 0.6765 0.4821
DACDR 0.1444 0.2278 0.2649 0.2891 0.3791 0.2675
PterRec 0.2187 0.3475 0.3961 0.4417 0.5599 0.3763
FDSA 0.3313 0.4667 0.5107 0.5644 0.6710 0.4840
Recformer 0.1332 0.2256 0.2685 0.2942 0.3987 0.2629
UniSRecID+t 0.2902 0.4328 0.4810 0.5361 0.6531 0.4513
IDP 0.3606 0.4966 0.5381 0.5935 0.6942 0.5105
IDPID+t 0.3566 0.4904 0.5363 0.5874 0.6987 0.5062
Impro. 1.10% 0.36% 0.35% 0.10% 0.80% 0.45%
@@ 4.8. Representing Intra-domain Cold-start Items via Pre-trained Item Embeddings (RQ5)
Dataset Metric HR@3 NDCG@3 HR@5 NDCG@5
Instruments IDPInner 0.1950 0.1655 0.2352 0.1819
Arts IDPInner 0.1294 0.1053 0.2705 0.1874
In this section, we aim to investigate the effects of generating IDs within the inner domain. This experimental result has the ability to demonstrate whether our IDP is capable of addressing inner-domain cold start issues. Instead of generating ID embeddings from the upstream domain to the downstream domain, we focus on generating ID embeddings within the inner domain. Specifically, for a domain $A$, we first train a sequential model. Then we use our IDP to generalize ID embedding for cold items (without any interaction). Last we directly evaluate the model’s performance on cold items without any tuning operation. Here, we randomly select 10% items as cold items and remove them from the training dataset. The results are shown in Table 8, where IDPInner is the model that re-generates ID embedding by our CDIM in the Inner domain. From this table, we can observe that our IDP can still provide reasonably accurate recommendations. This indicates that our IDP can be flexibly applied to inner-domain item cold start scenarios.
4.9. Extending to Multi-modal Scenario
Table 9. Results on of extending textual modality information to multi-modal information. The best performances are bold and underlined. All improvements are significant (t-test with p <{<}< 0.05).
Dataset Model NDCG@1 HR@3 NDCG@3 HR@5 NDCG@5 MRR
Arts SASRec 0.3228 0.4602 0.4027 0.5262 0.4298 0.4211
UniSRec 0.3841 0.5699 0.4926 0.6514 0.5261 0.5081
IDP 0.3404 0.4855 0.4249 0.5532 0.4528 0.4423
IDPID+t 0.4176 0.5849 0.5153 0.6581 0.5454 0.5302
IDPMM 0.3453 0.4860 0.4272 0.5493 0.4533 0.4436
[…]
Office SASRec 0.3408 0.5169 0.4429 0.6047 0.4790 0.4619
UniSRec 0.3899 0.5882 0.5055 0.6763 0.5418 0.5205
IDPID 0.3458 0.5227 0.4487 0.6088 0.4831 0.4668
IDPID+t 0.4096 0.5965 0.5186 0.6798 0.5528 0.5331
IDPMM 0.3461 0.5242 0.4497 0.6087 0.4845 0.4681
[…]
Instruments SASRec 0.3121 0.4562 0.3959 0.5252 0.4243 0.4149
UniSRec 0.3404 0.5274 0.4492 0.6135 0.4846 0.4686
IDPID 0.3286 0.4865 0.4207 0.5588 0.4504 0.4388
IDPID+t 0.3977 0.5715 0.4990 0.6481 0.5305 0.5149
IDPMM 0.3326 0.4867 0.4224 0.5568 0.4512 0.4401
[…]
Prime SASRec 0.1520 0.2743 0.2224 0.3503 0.2536 0.2584
UniSRec 0.1280 0.2518 0.1991 0.3302 0.2313 0.2372
IDPID 0.1539 0.2804 0.2268 0.3598 0.2595 0.2629
IDPID+t 0.1842 0.3209 0.2630 0.3996 0.2953 0.2965
IDPMM 0.1522 0.2825 0.2273 0.3616 0.2598 0.2628
[…]
In this paper, we aim to utilize the modality information to train our CDIM and transfer the ID embedding to downstream domains. In the Sec. 3.4 we mainly utilize the textual modality information as a bridge as it is more commonly available in the real world. However, our IDP can easily be extended to multi-modal scenarios. To evaluate it, we both utilize the image information and textual information to train the CDIM. Specifically, we first utilize pre-trained VIT(Dosovitskiy et al., 2020) to extract the image feature $im_{v}$ of items. Then we fuse the encoded textual feature $t_{v}$ and image feature $m_{v}$ by a two-layer MLP with a dropout layer, which is represented as $mm_{v}$. Subsequently, we replace the textual representation $t_{v}$ in Eq.5 and Eq.6 with multi-modal representation $mm_{v}$, and conduct the modality information contrastive learning (i.e., Eq.5) and behavior involved tuning (i.e., Eq.6). After that, we get a multi-modal-based CDIM, which can be used to transfer pre-trained ID embedding downstream.
[…]
We show the results in the Table.9. From this table, we have the following observations: (1) All variants of our IDP outperform SASRec and all various of our textual-enhanced IDP outperform UniSRec, which further demonstrates the power of our IDP. (2) By comparing IDP and IDPMM, we found that introducing multimodal information to train the CDIM module indeed helps improve the performance of our IDP. (3) By further comparing C and D, we found that in the case of text augmentation, C generally outperforms D. This further indicates the benefit of introducing multimodal information to our IDP. Furthermore, This experiment demonstrates the versatility of our IDP, which can be applied in multi-modal scenarios.
@@ 4.10. Analysis of Sequential Encoder (RQ6)
Dataset Metric HR@3 NDCG@3 HR@5 NDCG@5
Instruments IDP 0.4865 0.4207 0.5588 0.4504
IDP w/o PS 0.4843 0.4220 0.5522 0.4500
Prime IDP 0.2804 0.2268 0.3598 0.2595
IDP w/o PS 2912 0.2364 0.3715 0.2693
Dataset Metric HR@3 NDCG@3 HR@5 NDCG@5
Instruments SASRec 0.4562 0.3959 0.5252 0.4243
SASRecEX 0.4191 0.3553 0.4903 0.3846
Prime SASRec 0.2743 0.2224 0.3503 0.2536
SASRecEX 0.2259 0.1810 0.2946 0.2092